# Neutrino reheating predictions with non-thermal leptogenesis

### v1

@@ 1 Introduction
There is growing evidence that the early Universe undergoes an exponentially accelerated expansion era, known as inflation, which addresses both the flatness and horizon problems[2]. Inflation also seeds the anisotropy observed in the cosmic microwave background (CMB) and baryon acoustic oscillations (BAO) through quantum fluctuations. For reviews, see, e.g., refs. [3, 4]. In its simplest realization, the single-field slow-roll models, this accelerated expansion is driven by the potential energy of the inflaton, which rolls slowly towards the potential minimum. Inflation models are tested against the CMB anisotropy measurements, specifically the spectral tilt $n_{s}$ of the scalar power spectrum and the tensor-to-scalar ratio $r$. The Planck 2018 release determines the $n_{s}$ and $r$ at the pivot scale $k=0.002~{}\mathrm{Mpc}^{-1}$ to be [5]
$n_{s}=0.9649\pm 0.0126,\quad r<0.056,~{}(95\%~{}\mathrm{C.L.})\;.$ (1)
@@
In this work, we continue our investigation of the minimal framework by considering the reheating constraints on inflation models in detail. Assuming that inflaton couples only directly to RHNs, we introduce the concept of “neutrino reheating”, wherein the Universe is heated by neutrino-related interactions. In ref. [1], the inflationary observables are analyzed under generic assumptions regarding reheating, considering $50$ to $60$ e-folds from the end of inflation to horizon exit. In this work, we conduct a consistent calculation of inflationary observables within the context of neutrino reheating. We will demonstrate that neutrino reheating establishes a connection between the BAU and inflationary observables, providing specific predictions that may help break the degeneracy of inflationary models. Our focus is on the perturbative decay of the inflaton to RHNs, as Pauli blocking renders parametric resonance inefficient [39].
This work is organized as follows: in section 2 we review the reheating parameters and introduce the $N_{\mathrm{RH}}$-$N_{k}$ relation; in section 3 we introduce neutrino reheating from inflaton decay and calculate e-folds during neutrino reheating in two scenarios; in section 4, we examine four concrete inflationary models with neutrino reheating and calculate the corresponding parameter space; we conclude in section 5.
@@ 2 Reheating parameters and the $N_{\mathrm{RH}}$-$N_{k}$ relation
There are three quantities used to parameterize the unknown reheating dynamics: the effective equation of state parameter $\omega_{\mathrm{RH}}$, the duration of reheating in terms of e-folds $N_{\mathrm{RH}}$, and the reheating temperature $T_{\mathrm{RH}}$, which marks the temperature at the end of reheating. During reheating, the energy density can be expressed as $\rho=\rho_{\mathrm{end}}(a/a_{\mathrm{end}})^{-3(1+\omega_{\mathrm{RH}})}$, where $a$ is the scale factor and the subscripts end mark the end of inflation. By definition, we have
@@
where we use the energy density redshift relation. The energy density at the end of inflation is $\rho_{\mathrm{end}}=3V_{\mathrm{end}}/2$ (with the equation of state parameter $\omega=-1/3$), and the energy density at the end of reheating is $\rho_{\mathrm{RH}}=g_{*}\pi^{2}T_{\mathrm{RH}}^{4}/30$. Since the reheating temperature is unknown in many scenarios, one can re-express $T_{\mathrm{RH}}$ with $N_{\mathrm{RH}}$ and other quantities. The entropy conservation after reheating gives
@@
with $T_{\nu 0}=(4/11)^{1/3}T_{0}$. Then one finds
@@
Note that the same equation (10) can be obtained if $N_{\mathrm{RH}}=0$. $N_{\mathrm{RH}}=0$ means there is no separate reheating phase. Actually, $\omega_{\mathrm{RH}}=1/3$ is difficult to conceive since it corresponds to the equation of state of a radiation-dominated phase. The $N_{\mathrm{RH}}$-$N_{k}$ relation in (8) (and (9)) explicitly shows that reheating dynamics affects inflationary observables and is used to refine CMB constraints on inflation in refs. [8, 9, 10, 11, 40, 41].
There are three quantities related to inflationary models appearing in the $N_{\mathrm{RH}}$-$N_{k}$ relation, the potential energy at the end of inflation $V_{\mathrm{end}}$, the e-folds from the end of inflation to horizon exit $N_{k}$ and the Hubble parameter at horizon exit $H_{k}$. Once we specify a inflationary model, these quantities can be expressed in terms of the model parameters and then connected to reheating dynamics via eq. (8).
@@ 3 Neutrino reheating and non-thermal leptogenesis
where $c_{\mathrm{sph}}=28/79$ is the B--L to B conversion factor in sphaleron process, $\epsilon$ is the RHN decay asymmetry. Since the RHNs decay very quickly into radiation, overall, this scenario resembles the inflaton passing its energy directly to radiation, indicating $N_{\mathrm{RH}}=0$. It can be inferred from the two lower evolution plots in figures.1 and 2 in ref. [1]. As mentioned, $N_{\mathrm{RH}}=0$ also leads the $N_{\mathrm{RH}}$-$N_{k}$ relation to its reduced form (10). Meanwhile, from eq. (3), $N_{\mathrm{RH}}=0$ is equivalent to
@@
Since there exists a RHN matter-dominated phase, the reheating duration $N_{\mathrm{RH}}$ is nonzero. It can be calculated as follows. We first consider the possibility that the RHNs are produced relativistically with energy $E_{N}=M_{\phi}/2$ at $T_{\phi}$, where $T_{\phi}$ marks the time the inflaton decays efficiently. As the Universe cools down, the RHNs become non-relativistic at a temperature $T_{\mathrm{NR}}=T_{\phi}M_{N}/E_{N}$ and start matter-dominates the Universe. At $T_{\mathrm{RH}}$, RHNs efficiently decay to radiation, and the reheating ends. We can write the energy density at each temperature (time) [^2] Temperature may not be well defined without a thermal bath. $T_{\phi}$ and $T_{\mathrm{NR}}$ here should be understood symbolically as a measure of the cosmic time previous to (the cosmic time corresponds to) $T_{\mathrm{RH}}$.
@@
where $\rho_{N}|_{T_{\phi}}$ is $\rho_{N}$ at $T_{\phi}$. When the RHNs are relativistic, they redshift as radiation and feature an equation of state parameter $\omega_{\mathrm{I}}=1/3$. When RHNs become non-relativistic, they redshift as matter with an equation of state parameter $\omega_{\mathrm{II}}=0$. Since the main constituent in both phases is the RHNs, we can calculate the e-folds in these two phases as follows
$N_{\mathrm{I}}$ $=\frac{1}{3(1+\omega_{\mathrm{I}})}\mathrm{ln}\frac{\rho_{N}|_{T_{\phi}}}{\rho_{N}|_{T_{\mathrm{NR}}}}=\mathrm{ln}\frac{E_{N}}{M_{N}}\,.$ (18)
@@
$=\frac{1}{3}\left(\mathrm{ln}\frac{3}{2}V_{\mathrm{end}}\left(\frac{M_{N}}{E_{N}}\right)^{4}-\mathrm{ln}\frac{\pi^{2}g_{*}}{30}T_{\mathrm{RH}}^{4}\right)\,,$ (19)
where we first translate $\rho_{N}|_{T_{\mathrm{NR}}}$ to $\rho_{N}|_{T_{\phi}}$, then to $\rho_{\phi}$ using eq. (15). We also approximate $\rho_{\phi}|_{T_{\phi}}\simeq\rho_{\phi}|_{\mathrm{end}}$ by assuming that the inflaton decays soon after the end of inflation. Now we can easily express the reheating e-folds as $N_{\mathrm{RH}}=N_{\mathrm{I}}+N_{\mathrm{II}}$. From the expression of $N_{\mathrm{RH}}$, we can deduce the averaged equation of state parameter $\omega_{\mathrm{RH}}$ from eq. (3).
Now we are ready to discuss some limits of the parameters in the scenario we considered. First, all the e-folds should be non-negative. $N_{\mathrm{I}}$ in (18) naturally satisfies $N_{\mathrm{I}}\geq 0$, while $N_{\mathrm{II}}\geq 0$ sets an upper limit for $T_{\mathrm{RH}}$
$T_{\mathrm{RH}}\leq\left(\frac{30\rho_{N}|_{T_{\phi}}}{\pi^{2}g_{*}}\frac{M_{N}}{M_{\phi}/2}\right)^{1/4}\,.$ (20)
Second, since $\omega_{\mathrm{I}}=1/3$ and $\omega_{\mathrm{II}}=0$, the average reheating equation of state parameter is expected to be between 00 and $1/3$ and thus positive. As such, the $N_{\mathrm{RH}}$-$N_{k}$ relation in (9) indicates
$-\mathrm{ln}\frac{V_{\mathrm{end}}^{1/4}}{H_{k}}-\frac{1}{4}\mathrm{ln}\frac{45}{\pi^{2}g_{*}}+\frac{1}{3}\mathrm{ln}\frac{43}{11g_{*s,\mathrm{RH}}}-\mathrm{ln}\frac{k}{a_{0}T_{0}}-N_{k}>0\,.$ (21)
The equations (20) and (21) will help us locate the reasonable parameter space once the inflation model is specified.
In an intermediate scenario where the conditions for the limits do not apply, there is no analytical approximation for the final baryon asymmetry. One must solve the Boltzmann equations to understand the dynamics. From the numerical results of the Boltzmann equations, information can be obtained about the duration of the reheating phase $N_{\mathrm{RH}}$. Combined with the energy density at the end of inflation and at the end of reheating, the effective equation of state parameter $\omega_{\mathrm{RH}}$ can be deduced. With this information, predictions for neutrino reheating can be made using the $N_{\mathrm{RH}}$-$N_{k}$ relation.
The discussions on neutrino reheating presented above are general and apply to any inflation model. To obtain specific predictions, one must work with particular inflation models. We will present several examples in the next section.
@@ 4.1 Example 1: A polynomial type potential
The effects of reheating in this potential are discussed in refs. [42, 9, 10, 11]. We will show predictions in neutrino reheating together with non-thermal leptogenesis from inflaton decay.
Now we calculate the inflationary model-dependent quantities shown in the relation (8). The number of e-folds $N_{k}$ is
@@
where in the last equality, we omit the $\phi_{\mathrm{end}}$ term by assuming that the potential is flat enough such that $\phi_{k}\gg\phi_{\mathrm{end}}$. This relation allows us to express the slow-roll parameters (hence the spectral index and the tensor-to-scalar ratio) in terms of $N_{k}$.
@@
$H_{k}=\pi M_{\mathrm{P}}\sqrt{8A_{s}\epsilon_{k}}=2\pi M_{\mathrm{P}}\sqrt{\frac{\alpha}{\alpha+2}(1-n_{s})A_{s}}\,.$ (29)
@@
$V_{\mathrm{end}}=V_{k}\left(\frac{\phi_{\mathrm{end}}}{\phi_{k}}\right)^{\alpha}=3M_{\mathrm{P}}^{2}H_{k}^{2}\left(\frac{\phi_{\mathrm{end}}}{\phi_{k}}\right)^{\alpha}=12\pi^{2}M_{\mathrm{P}}^{4}\frac{\alpha}{\alpha+2}(1-n_{s})A_{s}\left(\frac{\alpha(1-n_{s})}{2(\alpha+2)}\right)^{\alpha/2}\,.$ (30)
Now, the three model-dependent quantities shown in (8) are all expressed in terms of the model parameter $\alpha$ and the spectral index $n_{s}$.
We need a non-zero inflaton mass to allow decaying to RHNs. Since $M_{\phi}^{2}=V^{{}^{\prime\prime}}|_{\mathrm{min}}$, it is nonzero only for $\alpha=2$ and we will focus on this case from now on. In this case, we have $M_{\phi}=m\simeq 8\times 10^{13}$ GeV, where the model parameter $m$ is calculated by matching the potential (22) at horizon exit with observations using the slow-roll approximation $V_{k}\simeq 3M_{\mathrm{P}}^{2}H_{k}^{2}$ with $H_{k}$ in (29).
@@
We first consider neutrino reheating in the instantaneous RHN decay limit, that is, $\Gamma_{N}\gg\Gamma_{\phi}$. Since $N_{\mathrm{RH}}=0$, we use the reduced $N_{\mathrm{RH}}$-$N_{k}$ relation (10) to find a prediction of $n_{s}=0.9652$. $N_{\mathrm{RH}}=0$ also leads to $\mathrm{ln}(3V_{\mathrm{end}}/2/\rho_{\mathrm{RH}})=0$ (13). Since $\rho_{\mathrm{RH}}=(\pi^{2}g_{*}/30)T_{\mathrm{RH}}^{4}$, one can express the reheating temperature $T_{\mathrm{RH}}$ as a function of $n_{s}$. The predicted $n_{s}$ corresponds to $T_{\mathrm{RH}}=2.79\times 10^{15}$ GeV. The final baryon asymmetry can be estimated from eq. (12) as [^3] We use $\epsilon=10^{-6}$ as a reasonable approximation of the RHN decay asymmetry. $Y_{\mathrm{B}}=1.85\times 10^{-5}$, several magnitudes larger than observed. The tensor-to-scalar ratio in this model (27) is a function of $\alpha$ and $n_{s}$. With $\alpha=2$ and the predicted $n_{s}$, we find $r=0.139$, which is not allowed given the latest constraints [6].
@@
Since nonzero inflaton mass requires $\alpha=2$, $N_{\mathrm{RH}}$ calculated from (18) and (19) depends on four parameters:
$M_{\phi},~{}M_{N},~{}T_{\mathrm{RH}},~{}n_{s}.$
The inflaton mass is calculated to be around $M_{\phi}\simeq 8\times 10^{13}$ GeV. We are left with three variables $M_{N},~{}T_{\mathrm{RH}},~{}n_{s}$. We choose two benchmark values of the RHN mass: $M_{N}=10^{11},10^{9}$ GeV. The $N_{\mathrm{RH}}$-$N_{k}$ relation allows to determine the parameter space in terms of $T_{\mathrm{RH}}$ and $n_{s}$. Using eq. (14), we can establish a relationship between the baryon asymmetry and the spectral index.
@@
Figure 1: The e-folds during reheating $N_{\mathrm{RH}}$, the reheating temperature $T_{\mathrm{RH}}$, the e-folds from the end of inflation to the exit of horizon $N_{k}$, the effective equation of state parameter during reheating $\omega_{\mathrm{RH}}$ in polynomial type inflation model ($\alpha=2$) with neutrino reheating in $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. The dark gray shaded region in the first column marks the Big Bang nucleosynthesis (BBN) forbidden region, and the light gray region is below the electroweak scale. The vertical light blue bands show allowed $1\sigma$ region of $n_{s}$ given by Planck 2018 [5]. The blue band marks $1\sigma$ region of a future CMB experiment [43, 44].
Figure 2: Predictions in polynomial type inflation model ($\alpha=2$) with neutrino reheating in $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. The dark blue line is the tensor-to-scalar ratio and the red dotted line is $Y_{\mathrm{B}}$. The shaded region show combined Planck [5] and BICEP/Keck [6] allowed region [42, 7]. The dashed light blue lines mark the $3\sigma$ region of $Y_{\mathrm{B}}$ given by Planck 2018 [12].

### v3

@@ 1 Introduction
There is growing evidence that the early universe undergoes an exponentially accelerated expansion era, known as inflation, which addresses both the flatness and horizon problems[2]. Inflation also seeds the anisotropy observed in the cosmic microwave background (CMB) and baryon acoustic oscillations through quantum fluctuations. For reviews, see, e.g., refs. [3, 4]. In its simplest realization, the single-field slow-roll models, this accelerated expansion is driven by the potential energy of the inflaton, which rolls slowly towards the potential minimum. Inflation models are tested against the CMB anisotropy measurements, specifically the spectral tilt $n_{s}$ of the scalar power spectrum and the tensor-to-scalar ratio $r$. The Planck 2018 release determines the $n_{s}$ and $r$ at the pivot scale $k=0.05~{}\mathrm{Mpc}^{-1}$ to be [5]
$n_{s}=0.9649\pm 0.0084,\quad r<0.063,~{}(95\%~{}\mathrm{C.L.})\;.$ (1)
@@
In this work, we continue our investigation of the minimal framework by considering the reheating constraints on inflation models in detail. Assuming that inflaton couples only directly to RHNs, we introduce the concept of “neutrino reheating”, wherein the universe is heated by neutrino-related interactions. In ref. [1], the inflationary observables are analyzed under generic assumptions regarding reheating, considering $50$ to $60$ e-folds from the end of inflation to horizon exit. In this work, we conduct a consistent calculation of inflationary observables within the context of neutrino reheating. We will demonstrate that neutrino reheating establishes a connection between the BAU and inflationary observables, providing specific predictions that may help break the degeneracy of inflationary models. Our focus is on the perturbative decay of the inflaton to RHNs, as Pauli blocking renders parametric resonance inefficient [39]. RHNs being responsible for reheating is also investigated in [40] and in the context of gravitational neutrino reheating [41, 42].
This work is organized as follows: in section 2 we review the reheating parameters and introduce the $N_{\mathrm{RH}}$-$N_{k}$ relation; in section 3 we introduce neutrino reheating from inflaton decay and calculate e-folds during neutrino reheating in two scenarios; in section 4, we analyze four concrete inflationary models with neutrino reheating and explore corresponding parameter spaces; we conclude in section 5.
@@ 2 Reheating parameters and the $N_{\mathrm{RH}}$-$N_{k}$ relation
Three quantities are used to parameterize the unknown reheating dynamics: the effective equation of state parameter $\omega_{\mathrm{RH}}$, the duration of reheating in terms of e-folds $N_{\mathrm{RH}}$, and the reheating temperature $T_{\mathrm{RH}}$, which marks the temperature at the end of reheating. During reheating, the energy density can be expressed as $\rho=\rho_{\mathrm{end}}(a/a_{\mathrm{end}})^{-3(1+\omega_{\mathrm{RH}})}$, where $a$ is the scale factor and the subscripts end mark the end of inflation. By definition, we have
@@
where in the last equality we use the energy density redshift relation. The energy density at the end of inflation is $\rho_{\mathrm{end}}=3V_{\mathrm{end}}/2$ (with the equation of state parameter $\omega=-1/3$), and the energy density at the end of reheating is $\rho_{\mathrm{RH}}=g_{*}\pi^{2}T_{\mathrm{RH}}^{4}/30$. Since the reheating temperature is unknown in many scenarios, one can re-express $T_{\mathrm{RH}}$ with $N_{\mathrm{RH}}$ and other quantities. The entropy conservation after reheating gives
@@
where $T_{\nu 0}=(4/11)^{1/3}T_{0}$. Then one finds
@@
Note that the same equation (10) can be obtained if $N_{\mathrm{RH}}=0$. $N_{\mathrm{RH}}=0$ means there is no separate reheating phase. Actually, $\omega_{\mathrm{RH}}=1/3$ is difficult to conceive since it corresponds to the equation of state of a radiation-dominated phase. The $N_{\mathrm{RH}}$-$N_{k}$ relation in (8) (and (9)) explicitly shows that reheating dynamics affects inflationary observables and is used to refine CMB constraints on inflation in refs. [8, 9, 10, 11, 43, 44, 45, 46, 47, 48].
There are three quantities related to inflationary models appearing in the $N_{\mathrm{RH}}$-$N_{k}$ relation, the potential energy at the end of inflation $V_{\mathrm{end}}$, the e-folds from the end of inflation to the horizon exit $N_{k}$ and the Hubble parameter at the horizon exit $H_{k}$. Once we specify an inflationary model, these quantities can be expressed in terms of the model parameters and then connected to reheating dynamics via eq. (8).
@@ 3 Neutrino reheating and non-thermal leptogenesis
where $c_{\mathrm{sph}}=28/79$ is the B--L to B conversion factor in sphaleron processes, $\epsilon$ is the RHN decay asymmetry. Since the RHNs decay very quickly into radiation, overall, this scenario resembles the inflaton passing its energy directly to radiation, indicating $N_{\mathrm{RH}}=0$. It can be inferred from the two lower evolution plots in figures.1 and 2 in ref. [1]. As mentioned, $N_{\mathrm{RH}}=0$ also leads the $N_{\mathrm{RH}}$-$N_{k}$ relation to its reduced form (10). Meanwhile, from eq. (3), $N_{\mathrm{RH}}=0$ is equivalent to
@@
Since there exists an RHN matter-dominated phase, the reheating duration $N_{\mathrm{RH}}$ is nonzero. It can be calculated as follows. We first consider the possibility that the RHNs are produced relativistically with energy $E_{N}=M_{\phi}/2$ at $T_{\phi}$, where $T_{\phi}$ marks the time the inflaton decays efficiently. As the universe cools down, the RHNs become non-relativistic at a temperature $T_{\mathrm{NR}}=T_{\phi}M_{N}/E_{N}$ and begin to dominate the universe as matter. At $T_{\mathrm{RH}}$, RHNs decay efficiently to radiation, and the reheating ends. We can write the energy density at each temperature (time) [^2] Note that $T_{\mathrm{RH}}$ may not be the highest temperature during reheating [49]. Additionally, temperature may not be well defined without a thermal bath. $T_{\phi}$ and $T_{\mathrm{NR}}$ here should be understood symbolically as a measure of the cosmic time previous to (the cosmic time corresponds to) $T_{\mathrm{RH}}$.
@@
where $\rho_{N}|_{T_{\phi}}$ is $\rho_{N}$ at $T_{\phi}$. When the RHNs are relativistic, they redshift as radiation and feature an equation of state parameter $\omega_{\mathrm{I}}=1/3$. When the RHNs become non-relativistic, they redshift as matter with an equation of state parameter $\omega_{\mathrm{II}}=0$. Since the main constituent in both phases is the RHNs, we can calculate the e-folds in these two phases as follows
$N_{\mathrm{I}}$ $=\frac{1}{3(1+\omega_{\mathrm{I}})}\mathrm{ln}\frac{\rho_{N}|_{T_{\phi}}}{\rho_{N}|_{T_{\mathrm{NR}}}}=\mathrm{ln}\frac{M_{\phi}}{2M_{N}}\,.$ (18)
@@
$=\frac{1}{3}\left(\mathrm{ln}\frac{720}{\pi^{2}g_{*}}+\mathrm{ln}\frac{V_{\mathrm{end}}}{M_{\phi}^{4}}+4\mathrm{ln}\frac{M_{N}}{T_{\mathrm{RH}}}\right)\,,$ (19)
where we first translate $\rho_{N}|_{T_{\mathrm{NR}}}$ to $\rho_{N}|_{T_{\phi}}$, then to $\rho_{\phi}|_{T_{\phi}}$ using eq. (15). We also approximate $\rho_{\phi}|_{T_{\phi}}\simeq\rho_{\phi}|_{\mathrm{end}}$ by assuming that the inflaton decays soon after the end of inflation. Now we can easily express the reheating e-folds as $N_{\mathrm{RH}}=N_{\mathrm{I}}+N_{\mathrm{II}}$. From the expression of $N_{\mathrm{RH}}$, we can deduce the averaged equation of state parameter $\omega_{\mathrm{RH}}$ from eq. (3).
Now we are ready to discuss some constraints on the parameters in the scenario we considered. First, all the e-folds should be non-negative. $N_{\mathrm{I}}$ in (18) naturally satisfies $N_{\mathrm{I}}\geq 0$, while $N_{\mathrm{II}}\geq 0$ sets an upper limit for $T_{\mathrm{RH}}$
$T_{\mathrm{RH}}\leq\left(\frac{30\rho_{N}|_{T_{\phi}}}{\pi^{2}g_{*}}\right)^{1/4}\frac{2M_{N}}{M_{\phi}}\,.$ (20)
A lower limit on $T_{\mathrm{RH}}$ is set by the Big Bang nucleosynthesis (BBN) to be a few MeV [50, 51, 52]. Second, since $\omega_{\mathrm{I}}=1/3$ and $\omega_{\mathrm{II}}=0$, the average reheating equation of state parameter is expected to be between 00 and $1/3$ and thus positive. As such, the $N_{\mathrm{RH}}$-$N_{k}$ relation in (9) indicates
$-\mathrm{ln}\frac{V_{\mathrm{end}}^{1/4}}{H_{k}}-\frac{1}{4}\mathrm{ln}\frac{45}{\pi^{2}g_{*}}+\frac{1}{3}\mathrm{ln}\frac{43}{11g_{*s,\mathrm{RH}}}-\mathrm{ln}\frac{k}{a_{0}T_{0}}-N_{k}>0\,,$ (21)
which gives an upper limit on $n_{s}$. A lower limit on $n_{s}$ can be determined by imposing a lower limit on $N_{k}$, which ensures that inflation solves the flatness and horizon problems and that reheating occurs above the BBN scale [11]. The discussions above will help identify a reasonable parameter space once an inflation model is specified.
In an intermediate scenario where the analytical approximations for the final baryon asymmetry do not apply, one must solve the Boltzmann equations to understand the dynamics. From the numerical results of the Boltzmann equations, information can be obtained about the duration of the reheating phase $N_{\mathrm{RH}}$. Combined with the energy density at the end of inflation and at the end of reheating, the effective equation of state parameter $\omega_{\mathrm{RH}}$ can be deduced. With this information, predictions for neutrino reheating can be made using the $N_{\mathrm{RH}}$-$N_{k}$ relation.
The discussions on neutrino reheating presented above are general and apply to any inflation model. To obtain specific predictions, one must work with a particular inflation model. We will present several examples in the next section.
@@ 4.1 Example 1: A polynomial type potential
The effects of reheating in this potential are discussed in refs. [53, 9, 10, 11]. Now we calculate the inflationary model-dependent quantities shown in the $N_{\mathrm{RH}}$-$N_{k}$ relation (8). The number of e-folds $N_{k}$ is
@@
where $M_{\mathrm{P}}$ is the reduced Planck mass. In the last equality, we omit the $\phi_{\mathrm{end}}$ term by assuming that the potential is flat enough such that $\phi_{k}\gg\phi_{\mathrm{end}}$. This relation allows us to express the slow-roll parameters (hence the spectral index and the tensor-to-scalar ratio) in terms of $N_{k}$.
@@
$H_{k}=\pi M_{\mathrm{P}}\sqrt{8A_{s}\epsilon_{k}}=2\pi M_{\mathrm{P}}\sqrt{\frac{\alpha}{\alpha+2}A_{s}(1-n_{s})}\,.$ (29)
@@
$V_{\mathrm{end}}=V_{k}\left(\frac{\phi_{\mathrm{end}}}{\phi_{k}}\right)^{\alpha}=3M_{\mathrm{P}}^{2}H_{k}^{2}\left(\frac{\phi_{\mathrm{end}}}{\phi_{k}}\right)^{\alpha}=12\pi^{2}M_{\mathrm{P}}^{4}\frac{\alpha}{\alpha+2}A_{s}(1-n_{s})\left(\frac{\alpha(1-n_{s})}{2(\alpha+2)}\right)^{\alpha/2}\,.$ (30)
Now, the three model-dependent quantities shown in the $N_{\mathrm{RH}}$-$N_{k}$ relation in eq. (8) are all expressed in terms of the model parameter $\alpha$ and the spectral index $n_{s}$.
We need a non-zero inflaton mass to allow decaying to RHNs. Since $M_{\phi}^{2}=V^{{}^{\prime\prime}}|_{\mathrm{min}}$, it is nonzero only for $\alpha=2$ and we will focus on this case from now on. In this case, we have $M_{\phi}=m\simeq 8\times 10^{13}$ GeV, where the model parameter $m$ is calculated by matching the potential (22) at horizon exit with observations using the slow-roll approximation $V_{k}\simeq 3M_{\mathrm{P}}^{2}H_{k}^{2}$ with $H_{k}$ in eq. (29).
@@
We first consider neutrino reheating in the instantaneous RHN decay limit, that is, $\Gamma_{N}\gg\Gamma_{\phi}$. Since $N_{\mathrm{RH}}=0$, we use the reduced $N_{\mathrm{RH}}$-$N_{k}$ relation (10) to find a prediction of $n_{s}=0.9652$. $N_{\mathrm{RH}}=0$ also leads to eq. (13). Since $\rho_{\mathrm{RH}}=(\pi^{2}g_{*}/30)T_{\mathrm{RH}}^{4}$, one can express the reheating temperature $T_{\mathrm{RH}}$ as a function of $n_{s}$ with relation determined by eq. (13). The predicted $n_{s}$ corresponds to $T_{\mathrm{RH}}=2.79\times 10^{15}$ GeV. The final baryon asymmetry can be estimated from eq. (12) as [^3] We use $\epsilon=10^{-6}$ as a reasonable approximation of the RHN decay asymmetry. $Y_{\mathrm{B}}=1.85\times 10^{-5}$, several orders of magnitude larger than observed. The tensor-to-scalar ratio in this model is a function of $\alpha$ and $n_{s}$ (27). With $\alpha=2$ and the predicted $n_{s}$, we find $r=0.139$, which is not allowed given the latest constraints [6].
@@
The neutrino reheating duration $N_{\mathrm{RH}}$ calculated from (18) and (19) depends on the following four parameters (since nonzero inflaton mass fixes $\alpha=2$): $M_{\phi},~{}M_{N}$, $M_{N}$, $T_{\mathrm{RH}}$ and $n_{s}$. Since the inflaton mass is calculated to be $M_{\phi}\simeq 8\times 10^{13}$ GeV, we are left with three variables
$M_{N},~{}T_{\mathrm{RH}},~{}n_{s}\,.$
@@
Figure 1: The e-folds during reheating $N_{\mathrm{RH}}$, the reheating temperature $T_{\mathrm{RH}}$, the e-folds from the end of inflation to the exit of horizon $N_{k}$, the effective equation of state parameter during reheating $\omega_{\mathrm{RH}}$ in polynomial type inflation model ($\alpha=2$) with neutrino reheating in $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. The dark gray shaded region in the first column marks the BBN forbidden region, and the light gray region is below the electroweak scale. The vertical light blue bands show allowed $1\sigma$ region of $n_{s}$ given by Planck 2018 [5]. The blue band marks $1\sigma$ region of a future CMB experiment [54, 55].
Figure 2: The full parameter space in polynomial type inflation model ($\alpha=2$) with neutrino reheating in the $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. In the left panel, the viable parameter space corresponds to the upper-left region above the red line. The dark gray shaded area at the top is excluded as it results in a negative $N_{\mathrm{I}}$. The vertical dashed line indicates the upper bound on $M_{N}$, derived from $M_{\phi}/2$. The lower limit on $M_{N}$ is adopted from [1]. In the right panel, the red curve shows the predicted $Y_{\mathrm{B}}$ as a function of $n_{s}$, while the green dotted line marks the observed baryon asymmetry value. The remaining shaded regions follow the same conventions as in the previous figure.
Figure 3: Predictions for the polynomial type inflation model ($\alpha=2$) with neutrino reheating in the $\Gamma_{N}\ll\Gamma_{\phi}$ scenario. The red dotted line shows the predicted tensor-to-scalar ratio, with the continuous part corresponds to the range of $n_{s}$ values that yield a sufficiently large baryon asymmetry $Y_{\mathrm{B}}$. The Planck-BK18 contours adopted from the latest update of Ref.[53], with detailed derivations provided in [7], incorporating both Planck and BICEP/Keck data. The ACT-LB-BK18 contour is based on [56], using Atacama Cosmology Telescope (ACT) DR6 data combined with CMB lensing and baryon acoustic oscillation (BAO) measurements. The P-ACT-LB-BK18 contour additionally includes Planck data [56].